OpenStudy (anonymous):

solve e^ (2 x)-6 e ^(x) = 8 , for x, x>1

sam (.sam.):

Let y= e^x y^2 -6y -8 = 0 y=7.123 ,y=-1.123 e^x = 7.123 x= ln(7.123)

sam (.sam.):

Ignore -1.123 since \[x >1\]

OpenStudy (anonymous):

\[Let, e ^{x}=a ,so , e ^{2x} =a ^{2}\] \[a ^{2}-6a=8\] \[a=4, a=-2\] \[e ^{x}=-2,4\] \[e ^{x}=-2 , Not -Possible\ \[So, e ^{x}=4\] \[x=2\ln 2\]

OpenStudy (anonymous):

thanks much